1. Calculate the force needed to give a mass of 5kg an acceleration of
a) 2m/s 2 b) 22 km/h (convert to m/s)
2. A constant force changes the speed of a 100kg sprinter from 4.0m/s to 6.0 m/s in 1.0 s. Calculate
a) The acceleration of the sprinter during this time.
b) The magnitude of the force
3. If a car of mass 900kg travelling at 30m/s must stop in 6 seconds, calculate
a) deceleration
b) force applied by the brakes
4. A force of 12 N is used to move a box of mass 20kg along a road. If there is a constant frictional force of 4N opposing the motion, what will be the acceleartion of the box?
5. A toy rocket of mass 10kg rises vertically upwards. The gases it emits give it an upward thrust of 120N. Calculate
a) the unbalanced force (resultant force) acting on the rocket.
b) The acceleration of the rocket 4 marks
F
1
Expert's answer
2019-09-17T12:05:06-0400

1.


"F=ma"

a)


"F=(5)(2)=10\ N"

b)


"F=(5)(22)\frac{1000}{3600^2}=0.0085\ N=8.5\ mN"

2.

a)


"a=\frac{6-4}{1}=2\frac{m}{s^2}"

b)


"F=(100)(2)=200\ N"

3. a)

"a=\frac{30}{6}=5\frac{m}{s^2}"

b)


"F=(900)(5)=4500\ N"

4.


"a=\frac{F-F_{fr}}{m}=\frac{12-4}{20}=0.4\frac{m}{s^2}"

5.a)


"F=T-mg=120-(10)(9.8)=22\ N"

b)


"a=\frac{22}{10}=2.2\frac{m}{s^2}"